If the equation z+(3+i)z2-3z- (m+i)=0, where m is a real number, has at least one real root, then m can have the value equal to


Answer: 1 or 5 

Grade:11

2 Answers

Ravi
9 years ago
Let a,b,c be the roots.
If a is real, (b,c are conjugates)
a+b+c=3+i
abc=m+i
ab+bc+ca=-3

Use this information and solve for m.
........................................................................................................................................
mycroft holmes
9 years ago
Solution: